### v1

@@ 4.5 Robustness analysis
Here we present a theorem justifying the estimation of infinitesimal generators of functions that are themselves approximations of an unknown ground truth function (e.g., functions learned from data). In the theorem, $f$ plays the role of a ground truth function whose generators we wish to estimate, and $\hat{f}$ plays the role of our estimate of $f$.
Theorem 4.1 (Approximate annihilators of a function have flows that are approximate invariances).
@@
Let $\hat{\Psi}(t,z)$ be the flow associated with $\hat{X}$.
Then
@@
Proof 4.2.
To prove (19), we write
(21) $\hat{X}(f)=\hat{X}(f-\hat{f}+\hat{f})=\hat{X}(f-\hat{f})+\hat{X}(\hat{f})=\hat{X}(f-\hat{f}),$
where the second equality is by linearity of $\hat{X}$, and the third is by the fact that $\hat{X}(\hat{f})$ annihilates $\hat{f}$.
Taking norms on both sides and using linearity of $\hat{X}$ and then the Lipschitz property,
(22) $\|\hat{X}(f)\|_{\infty}\leq\lambda\epsilon.$
@@
Thus, it remains to upper bound $\left|\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}\right|$ for all $\tau$. We do this as follows: by the chain rule,
The second equality is by definition of the flow associated with a vector field. The third equality is by definition of the vector field applied to functions. Applying (19), we finally get
(25) $\left|\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}\right|\leq\lambda\epsilon.$
Plugging this into (23) completes the proof of the theorem.
We can interpret Theorem 4.1 as saying that a vector field that annihilates an approximation of the ground truth function $f$ has a flow that approximately serves as an invariance for $f$, where the degree of approximation depends on the smoothness of the discovered vector field and the parameter $t$ of the flow. Lipschitzness of the vector field $\hat{X}$ can be constrained by constraining it to have, e.g., polynomial coefficients with bounded degrees and monomial coefficients.
@@ 4.6 Limitations
Our methods are state of the art for continuous symmetry detection and enforcement, both for computational reasons and due to the fact that our methods can readily handle continuous symmetries which are not affine transformations and without the need to discretize continuous transformations. However, there are limitations to our methods. First, as we have mentioned previously, is the issue which arises when vector fields $X$ and $fX$ are both present in the search space. We have suggested that a symbolic approach may resolve this, though non-symbolic approaches such as ours must grapple with this issue. Our approach of searching for polynomials of lower order first partially resolves this issue. Another promising resolution is to restrict the search space of vector fields to particular types of symmetries, such as isometries, as described in section 4.3. However, this issue is not fully resolved.
Another limitation is that our method requires an assumption about the coefficient functions of the vector fields. Herein, we primarily use polynomial functions, and although polynomials are universal approximators [19], the number of polynomial terms for a given coefficient function can increase rapidly, particularly when the number of dimensions is high. We believe this limitation is best resolved, as with the previous limitation, by restricting the search space to particular types of symmetries. However, the vector fields in the restricted search space must be calculated if they are not already known.
Though our method is comparatively efficient, the reliance on the elbow curve to determine both the number of discovered symmetries and, when level set estimation is used, the number of coefficients in the level set, is a limitation. This limitation is of a more persistent nature than the previous limitations, appearing in analogous form as the dimension hyperparameter in manifold learning methods, or as the number of clusters in $k$-means clustering. It is possible that our method would be improved by another means of selecting this number, however.
We also mention the limitations inherited via the choice of optimization of equations (14) and (15). Cost function selection, optimization algorithm selection, hyperparameter selection, together with the potential for poor parameter initialization all constitute limitations to our method. We believe the optimization task is much more straight-forward than training many other symmetry detection models, however, as most other methods require fine-tuning neural networks.
Lastly, our notion of similarity of vector fields, while adequate for our experiments, is limited. Primarily, when a collection of vector fields is estimated and taken to constitute the infinitesimal generators for the symmetry group, it is not clear how to compare the estimated collection of vector fields to a ground truth collection. The similarity score also depends on the range of a given dataset, so that two vector fields may have differing similarity scores for different datasets.
@@ 5 Experiments
We offer four experiments to demonstrate our extension of previous work in the area. First, we learn the infinitesimal symmetries of a machine learning model given as a neural network: while this has been suggested as possible in principle [23], it was not explicitly accomplished. The second experiment enforces non-affine symmetry in model training, obtaining high test accuracy without the need of data augmentation. A trajectory prediction experiment follows, wherein we regularize model training using functions estimated to be invariant with respect to the Hamiltonian vector field of a given Hamiltonian. Our final experiment discovers infinitesimal isometries of a classification function, where the distance measure in the original feature space is induced by a quadratic feature map.
5.1 Symmetry Discovery for Neural Networks
In our first experiment, we demonstrate how a feedforward neural network can learn a rotationally symmetric function from 2D input data and how its learned derivatives can be used to detect the underlying symmetry. We generate $N=1000$ points $(x,y)$ uniformly from a unit disk by first sampling
$r=\sqrt{U},\quad\theta\sim\mathrm{Uniform}(0,2\pi),\quad U\sim\mathrm{Uniform}(0,1),$
and then converting to Cartesian by means of $x=r\cos(\theta)$ and $y=r\sin(\theta)$. Our regression target is $f(x,y)=x^{2}+y^{2}$, which is invariant under planar rotations about the origin. A ground-truth infinitesimal generator for this rotational symmetry is
$X=-y\partial_{x}+x\partial_{y}.$
We define a multi-layer perceptron $F:\mathbb{R}^{2}\to\mathbb{R}$ with two hidden layers of size 64 and train it for 5000 epochs using the Adam optimizer at a learning rate of 0.001. The final mean-square error (MSE) on the training set is on the order of $10^{-5}$.
To uncover the underlying rotational symmetry from the network’s derivatives, we apply equation 14 under the assumption that the coefficients of the estimated vector field $\hat{X}$ are affine. The gradient of $F$ is approximated using a finite-difference method. Using the L1 Loss, the (Riemannian) Autograd optimizer with learning rate $0.01$, and training for $5000$ epochs, we recover the following estimate of the symmetry of $F$:
$\hat{X}=(0.0012x-0.7075y+0.0002)\partial_{x}+(0.7067x-0.0058y-0.0005)\partial_{y},$
which obtains a similarity score of $0.99999$ with respect to the ground truth. This result shows that our method can estimate the infinitesimal generators of functions defined by MLPs.

### v2

@@ 5.4 Robustness analysis
Here we present a theorem justifying the estimation of infinitesimal generators of functions that are themselves approximations of an unknown ground truth function (e.g., machine learning functions learned from data). In the theorem, $f$ plays the role of a ground truth function whose generators we wish to estimate, and $\hat{f}$ plays the role of an approximation of $f$.
Theorem 5.1.
@@
Let $\hat{\Psi}(t,z)$ be the flow associated with $\hat{X}$. Then
@@
Proof 5.2.
[…]
@@
(16) $|f(\hat{\Psi}(t,z))-f(z)|=|f(\hat{\Psi}(t,z))-f(\hat{\Psi}(0,z))|=\left|\int_{0}^{t}\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}~d\tau\right|\leq t\cdot\sup_{\tau\in[0,t]}\left|\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}\right|.$
Thus, it remains to upper bound $\left|\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}\right|$ for all $\tau$. By the chain rule,
The second equality is by definition of the flow associated with a vector field. The third equality is by definition of the vector field applied to functions. Applying Equation (14), we finally get $\left|\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}\right|\leq\lambda\epsilon.$ Plugging this into Equation (16) completes the proof.
Theorem 5.1 says that a vector field that annihilates an approximation of the ground truth function $f$ has a flow that approximately serves as an invariance for $f$, where the degree of approximation depends on the smoothness of the discovered vector field and the parameter $t$ of the flow. Lipschitzness of the vector field $\hat{X}$ can be achieved by constraining it to have, e.g., polynomial coefficients with bounded degrees and monomial coefficients.
5.5 Comparison of Discovered Symmetries to Ground-Truth Symmetries
To evaluate symmetry discovery methods, it is important to quantify the ability of a given method to recover specific ground truth symmetries in synthetic experiments. This is a distinct problem from evaluating the extent to which the discovered symmetry is approximately an infinitesimal generator of an invariant flow, and thus evaluation via the empirical loss is not sufficient. For a ground truth vector field $X$ and an estimated vector field $\hat{X}=\sum_{i=1}^{N}\hat{f}_{i}\partial_{x^{i}}$, we define the angle between them as
@@ 5.5 Comparison of Discovered Symmetries to Ground-Truth Symmetries
where $\langle X,\hat{X}\rangle_{g}=\sum_{i,j}f_{i}\hat{f}_{j}g_{ij}$, $X||_{g}=\sqrt{\langle X,X\rangle_{g}}$, and where $\mathbb{E}\left[u(\mathbf{x})\right]=\int_{\Omega}u(\mathbf{x})d\mathcal{M},$ with the region $\Omega$ being defined by the range of a given dataset. Ordinarily, this is the full range of the dataset. This formula is a generalization of the formula given in [Shaw1] in the case where the manifold and/or metric is not assumed to be Euclidean.
The notion of a “cosine similarity” between vector fields also induces a method by which the extent to which a particular smooth function is invariant can be quantified in relation to other functions. For a fixed vector field $X$, a function $f$ is $X$-invariant if and only if $X$ is orthogonal to the gradient of $f$, which vector field we denote $X_{f}$. Thus, the extent to which $f$ is $X$-invariant can be quantified in terms of the cosine of the angle between $X$ and $X_{f}$, given in Equation (18): the closer to 0 this value is, the more $X$-invariant $f$ is.
5.6 Symmetry Enforcement
Here, we detail two approaches to enforcement of discovered symmetries in downstream machine learning models. In the first, we show that we can construct a polynomial basis for the space of features invariant to a given vector field $X$. By taking polynomial degrees to be sufficiently large, we can guarantee that sufficiently smooth machine learning models trained on these features are arbitrarily close to invariant to $X$. The second method induces approximate invariance in a new model by regularization of the training objective based on $X$.
5.6.1 Construction of a basis for the invariant feature space
In this section, we assume that we have already estimated a vector field $X$ that annihilates the target function $f$. Our next goal is to construct a set of feature functions that are (i) simple, (ii) by design are also annihilated by $X$ and thus share the same invariance as $f$, and (iii) can be combined to approximate a large family of “well behaved” functions annihilated by $X$, at least on some domain. Concretely, we seek the following properties for our invariant feature set:
  1. 1.
    Soundness: No machine learning model that is not $X$-invariant can be constructed from the invariant feature set.
  2. 2.
    Completeness: “Every” $X$-invariant functional can be approximated by some machine learning model built on the invariant features.
We do not claim to construct an invariant feature set which satisfies these properties exactly: rather, we seek to approximately satisfy these properties.
We note that $X$ is a linear differential operator and that the space $C^{\infty}(\Omega)$ (for an arbitrary compact set $\Omega\subseteq\mathbb{R}^{d}$) can be endowed with a norm-inducing inner product: $\left\langle v,w\right\rangle:=\int_{\Omega}v(z)w(z)~dz.$ With this inner product and norm in mind, we will produce a sequence of bases $B_{k}:=\{f_{1,k},...,f_{k,k}\}$ with spans $H_{k}$ and $C$-approximate spans $H_{k,C}$ satisfying the following:
Definition 5.3 (Completeness and soundness).
We say that a sequence $\{B_{k}\}_{k=1}^{\infty}$ is complete for $X$ if for all $F\in\ker(X)$ and $\epsilon>0$, there exists a sequence index $k_{0}$ such that for all $k\geq k_{0}$, $\min_{\hat{F}\in H_{k}}\|F-\hat{F}\|_{2}\leq\epsilon.$
We say that $\{B_{k}\}_{k\geq 1}$ is sound for $X$ if for all error tolerances $\epsilon>0$ and coefficient bounds $C>0$, there exists a sequence index $k_{0}$ such that for all $k\geq k_{0}$ and $F_{k}\in H_{k,C}$, $\|XF_{n}\|_{\infty}\leq\epsilon.$
Least squares approach to basis construction
Ultimately, we are trying to solve a first-order linear PDE. We thus adopt the intuition from linear least squares. We fix a basis $\hat{B}:=\{p_{j}\}_{j\geq 1}$ for the space of $C^{\infty}$ functions. For example, this could be a basis of orthonormal polynomials (but this is not needed for our construction). We write also $\hat{B}_{k}:=\{p_{j}\}_{j=1}^{k}$ and $\hat{H}_{k}:=\text{span}\{\hat{B}_{k}\}$. We also write $\text{span}_{L_{1},B}\{S\}$ for the set of linear combinations of elements of any set $S$ with coefficients $L_{1}$-bounded by $B>0$. We would like to solve $Xf=0$ for $f\in\hat{H}_{k}$, but in general only the trivial solution exists.
We perform the following procedure, which returns for any given $k\in\mathbb{N}$ a set of approximate basis functions $\{f_{j}\}_{j=1}^{k}$ for $\ker(X)\cap\hat{H}_{k}$:
  1. 1.
    Form the matrix $Z:=[Xp_{1},...,Xp_{k}].$
  2. 2.
    We look at the smallest singular values of the matrix $Z^{\dagger}Z$. These correspond to eigenvectors that are coefficient vectors in $\ker(Z)$. Note that we are viewing $Z$ as a linear operator from $\mathbb{R}^{k}\to C^{\infty}$. Then $\ker(Z)$ is the set of coefficient vectors generating linear combinations of elements of $\hat{B}$ that lie in $\ker(X)$.
    Specifically, let the singular values of $Z^{\dagger}Z$ be $\sigma_{1}\geq\cdots\geq\sigma_{k}$, and let $v_{1},...,v_{k}\in\mathbb{R}^{k}$ denote the corresponding singular vectors (eigenvectors). We denote by $k_{0}:=\min\{j\in[k]~|~\sigma_{j}\leq\epsilon\}$. Here, we recall that $\epsilon>0$ is an error tolerance parameter.
    Define the functions $\{f_{j}\}_{j=1}^{k-k_{0}-1}$ as follows: $f_{j}:=\sum_{\ell=1}^{k}v_{k_{0}+j-1,\ell}\cdot p_{\ell}.$
Having established a set of functions $\{f_{j}\}_{j=1}^{k-k_{0}-1}$, our next task is to establish completeness and soundness for this set, which we do in the Sections 5.6.1 and 5.6.1.
Establishing completeness
To establish completeness, we need to figure out a natural restriction of the class of invariant $C^{\infty}$ functions that can be uniformly approximated by the restricted span of $\{f_{j}\}_{j=1}^{k}$ by taking $k$ large enough. We consider the following approximation target class.
Definition 5.4 (Approximation target class).
For any $\gamma>0,k\in\mathbb{N},B\in(0,\infty]$, we denote by $F_{\gamma,k,B}$ the set of functions $g\in\ker(X)$ satisfying the following: $\|g-\text{span}_{L_{1},B}\{v_{1},...,v_{k}\}\|_{2}\leq\gamma.$ We call $F_{\gamma,k,B}$ the approximation target class.
The significance of the approximation target class is that every element of $\ker(X)$ lies in some approximation target class for large enough $k$ and $B$. This is because $\{p_{j}\}_{j=1}^{\infty}$ is a basis for the space of $C^{\infty}$ functions on $\Omega$.
Lemma 5.5 (Small $Xg$ in $\hat{H}_{k}$ implies closeness to $\text{span}\{f_{1},...,f_{k}\}$).
Let $\gamma>0$. Suppose that a function $g\in\hat{H}_{k}$ satisfies $\|Xg\|_{2}\leq\gamma.$ Then $\|g-\text{span}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}\leq\frac{\gamma}{\sqrt{\sigma_{k_{0}-1}}}\leq\frac{\gamma}{\sqrt{\epsilon}}.$
Proof 5.6.
We will use the fact that $\{v_{1},...,v_{k}\}$ is an orthonormal basis. We write
$\gamma^{2}$ $\geq\|Xg\|_{2}\geq\|Xg-X\text{span}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}^{2}=\sum_{j=1}^{k_{0}-1}c_{j}^{2}\cdot\|v_{j}\|_{2}^{2}\sigma_{j}\geq\sigma_{k_{0}-1}\cdot\sum_{j=1}^{k_{0}-1}c_{j}^{2}\|v_{j}\|_{2}^{2}$
$\geq\epsilon\|g-\text{span}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}^{2}.$
The first inequality is by the hypothesis of the theorem. The second inequality is using orthonormality of $\{v_{1},...,v_{k}\}$. Specifically, $\{v_{k_{0}},...,v_{k}\}$ is orthogonal to $\{v_{1},...,v_{k_{0}-1}\}$. The third inequality is by the fact that $\sigma_{i}\geq\sigma_{j}$ whenever $i\leq j$. Finally, the last inequality is by definition of $k_{0}$. This completes the proof.
Theorem 5.7 (Approximating the approximation target class via approximate kernels).
Let $\gamma>0$ be arbitrary. For any $k\in\mathbb{N}$ and $B>0$, for any $g\in F_{\gamma,k,B}$, we have
$\|g-\text{span}_{L_{1},B}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}\leq\gamma+\frac{\lambda_{X}\gamma}{\sqrt{\epsilon}}.$
Proof 5.8.
By definition of $F_{\gamma,k,B}$, there exists some $\hat{g}\in\text{span}_{L_{1},B}\{v_{1},...,v_{k}\}$ such that $\|g-\hat{g}\|_{2}\leq\gamma.$ Again from the definition of $F_{\gamma,k,B}$, we have that $Xg=0$. This implies that $\|X\hat{g}\|_{2}=\|X\hat{g}-Xg\|_{2}\leq\lambda_{X}\|\hat{g}-g\|_{2}\leq\lambda_{X}\cdot\gamma.$ Here, we have used the Lipschitz property of $X$, with $L_{2}\to L_{2}$ Lipschitz constant $\lambda_{X}$.
From Lemma 5.5, $\hat{g}$ is close to $\text{span}\{f_{1},...,f_{k-k_{0}-1}\}$: $\|\hat{g}-\text{span}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}\leq\frac{\lambda_{X}\gamma}{\sqrt{\epsilon}}.$ Since $\hat{g}$ has coefficients for $\{v_{1},...,v_{k}\}$ that are bounded by $B$ in the $L_{1}$ norm, this can be strengthened to $\|\hat{g}-\text{span}_{L_{1},B}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}\leq\frac{\lambda_{X}\gamma}{\sqrt{\epsilon}}.$
Putting everything together using the triangle inequality, we have that
$\|g-\text{span}_{L_{1},B}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}\leq\|g-\hat{g}\|_{2}+\|\hat{g}-\text{span}_{L_{1},B}\{f_{1},...,f_{k-k_{0}-1}\}\|_{2}\leq\gamma+\frac{\lambda_{X}\gamma}{\sqrt{\epsilon}}.$
Plugging in $\gamma=\epsilon$, we get that (5.7) converges to 0 as $\epsilon\to 0$.
Establishing soundness
To establish soundness, we restrict the span of the $\{f_{j}\}$ to have $B$-bounded coefficients (in the $L_{1}$ norm). This will allow us to upper bound $\|Xg\|_{2}$ for arbitrary $g$ in this set. This upper bound implies an upper bound on the $L_{1}$ norm that grows with the volume of the domain $\Omega$. This, in turn, implies that no non-negligible subset can have $Xg$ too large in $L_{\infty}$ norm.
Theorem 5.9 (Soundness).
Let $B>0$. We have that for any $g\in\text{span}_{L_{1},B}\{f_{1},...,f_{k-k_{0}-1}\}$, $\|Xg\|_{2}\leq\sqrt{B\epsilon}.$ This implies $\|Xg\|_{1}\leq\text{Vol}(\Omega)\cdot\|Xg\|_{2}\leq\text{Vol}(\Omega)\sqrt{B\epsilon}.$
Proof 5.10.
The first inequality follows immediately from the choice of $k_{0}$ and the $L_{1}$ bound $B$. The second inequality follows from elementary properties of norms.
We note that soundness has no dependence on $k$, because of the choice of $k_{0}$.
5.6.2 Enforcing symmetries via regularization
We now turn to the topic of enforcing continuous symmetries in a smooth machine learning model during training using regularization based on discovered infinitesimal generators $\{X_{k}\}_{k=1}^{s}$.
Suppose that we seek to learn a function $f$ mapping data instances $x_{i}$ to targets $y_{i}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{m}$. Suppose also that the function $f$ is estimated by means of the minimization of $L(f(\mathbf{x}),\mathbf{y})$ for a smooth loss function $L$.
Now suppose that we also desire that the function $f$ learns to exhibit continuous symmetry with respect to the independent infinitesimal generators $\{X_{k}\}_{k=1}^{s}$. That is, for each component $f_{j}$ of $f$, $X_{k}(f_{j})=0,$ for $1\leq k\leq s$ and $1\leq j\leq m$. Our method of symmetry enforcement in training is to minimize the following loss function:
The loss function given in Eq. (19) resembles that of a PINN, where our $\tilde{L}$ term corresponds to differential constraints imposed by the governing equations for a given physical system [Raissi1]. Additionally, our symmetry enforcement term is a differential constraint, so that our method of symmetry enforcement appears to be a special case of a PINN in which the differential operators are linear, first-order, and homogeneous. Despite the mathematical similarity of enforcing continuous symmetries using vector fields and the enforcement of differential equations governing a physical system, we note that PINNs do not erase the need for symmetry enforcement using vector fields. This is due simply to the fact that the desired symmetry characteristics of a machine learning model are not always readily expressible in terms of a physical system. We note, however, that a machine learning task subject to physical constraints may be amenable to vector field regularization–in particular, a Hamiltonian vector field can often be readily given for a Hamiltonian defined on a symplectic manifold.
@@ 5.7 Limitations
Our methods are state of the art for continuous symmetry detection and enforcement, both for computational reasons and due to the fact that our methods can readily handle continuous symmetries that are not affine transformations without the need to discretize continuous transformations. However, there are limitations to our methods. First, as we have mentioned previously, is the issue which arises when vector fields $X$ and $fX$ are both present in the search space. We have suggested that a symbolic approach may resolve this, though non-symbolic approaches such as ours must grapple with this issue. Our approach of restricting the model class to low-degree polynomials partially resolves this issue. Another promising resolution is to restrict the search space of vector fields to particular types of symmetries, such as isometries, as described in Section 5.2. Further work is required to fully resolve this.
Another limitation is that our method requires an assumption about the coefficient functions of the vector fields. Herein, we primarily use polynomial functions, and although polynomials are universal approximators [PolyApprox], the number of polynomial terms for a given coefficient function can increase rapidly, particularly when the number of dimensions is high. We believe this limitation is best resolved, as with the previous limitation, by restricting the search space to particular types of symmetries.
Though our method is comparatively efficient, the reliance on the elbow curve to determine both the number of discovered symmetries and, when level set estimation is used, the number of coefficients in the level set, is a limitation. Unlike the previous limitations, this limitation is shared by methods for other common problems, appearing in analogous form as the dimension hyperparameter in manifold learning methods, or as the number of clusters in $k$-means clustering. It is possible that our method could be improved by another means of selecting this number.
We also mention the limitations inherited via the choice of optimizing Equation (12) Cost function selection, optimization algorithm selection, hyperparameter selection, together with the potential for poor parameter initialization all constitute limitations to our method. However, the optimization task is much more straight-forward than training many other symmetry detection models, as most other methods require fine-tuning neural networks.
@@ 6 Experiments
We offer three experiments to provide for an effective demonstration of our methods. First, we learn the infinitesimal symmetries of a machine learning model given as a neural network. While this has been previously suggested as possible in principle (see [Shaw1]), it was not explicitly accomplished. The second experiment enforces non-affine symmetry in model training, obtaining high test accuracy without the need of data augmentation. Our final experiment discovers infinitesimal isometries of a classification function, where the distance measure in the original feature space is induced by a quadratic feature map.
6.1 Symmetry Discovery and Enforcement for Neural Networks
In our first experiment, we demonstrate how a feedforward neural network can learn a rotationally symmetric function from 2D input data,how its estimated derivatives can be used to detect the underlying symmetry, and how that learned symmetry can be used to further enforce model symmetry for better generalization. For the training set, we generate $N=1000$ points $(x,y)$ uniformly from a unit disk by first sampling $r=\sqrt{U},\quad\theta\sim\mathrm{Uniform}(0,\pi),\quad U\sim\mathrm{Uniform}(0,4),$ and then converting to Cartesian coordinates by means of $x=r\cos(\theta)$ and $y=r\sin(\theta)$. The test set is sampled similarly, but with $N=3000$ and $\theta\sim\mathrm{Uniform}(\pi,2\pi)$.
Our regression target is $f(x,y)=x^{2}+y^{2}$, which is invariant under planar rotations about the origin. A ground-truth infinitesimal generator for this rotational symmetry is $X=-y\partial_{x}+x\partial_{y}.$ We define a multi-layer perceptron $F:\mathbb{R}^{2}\to\mathbb{R}$ with two hidden layers of size 64 and train it for 5000 epochs using the Adam optimizer at a learning rate of $0.01$. The final mean-square error (MSE) on the training set is on the order of $10^{-5}$, while the test MSE is approximately $0.8181$.
To uncover the underlying rotational symmetry from the network’s derivatives, we apply Equation (12) under the assumption that the coefficients of the estimated vector field $\hat{X}$ are affine. The gradient of $F$ is approximated using autograd. Using the L1 Loss, the (Riemannian) Adagrad optimizer with learning rate $0.1$, and training for $1000$ epochs, we recover the following estimate of the symmetry of $F$: $\hat{X}=(0.0012x+0.7144y-0.0135)\partial_{x}+(-0.6996x-0.0059y+0.0037)\partial_{y},$
We now use the learned symmetry to retrain the model with symmetry regularization. We train a new model as before, only with $4500$ epochs instead of $5000$. Then, for the remaining $500$ epochs, we regularize with symmetry using equation 19 with $\lambda=0.01$, $\mathcal{L}_{2}$ being the MSE loss, and a new learning rate of $0.001$. This results in a test MSE (for labels) of $0.1148$, a fraction of the test MSE of $0.8181$ obtained without symmetry enforcement.